166 Preliminary Report

Operations Report

SITE 1007
(Proposed Site BT-3)

Transit: Site 1003 to Site 1007
The 4.2 nmi transit to Site 1007 was made in dynamic positioning mode while tripping pipe and handling the BHAs. A retrievable beacon was dropped at 0755 hr on the GPS coordinates for Site 1007 (24°30.26'N, 79°19.34'W).

Hole 1007A
An APC/XCB BHA, similar to that used at previous sites, was assembled and run to the seafloor. Hole 1007A was spudded at 1110 hr on 24 March. On Core 1007A-1H, the flapper core catcher remained open and some core fell out on the rig floor. The core had a good mudline and was archived as 6.65 m recovery; however, another mudline core was taken as a precaution. The water depth was 650.3 mbsl based on recovery of the mudline core.

Hole 1007B
The ship was not moved and Hole 1007B was initiated at 1135 hr on 24 March. Core 1H recovered 9.90 m of core and appeared to have a good mudline. The water depth was 647.4 mbsl based on recovery. APC Cores 1007B-1H to 10H were taken from 0 to 38.0 and 40.0 to 91.7 mbsf with 97.1% recovery. The APC core attempt after Core 4H was rejected at 38.0 mbsf. A 1-m-thick hard layer (38-39 mbsf) was penetrated by drilling 2 m (38.0-40.0 mbsf), and APC coring was resumed. Adara temperature tool heat-flow measurements were taken during coring of Cores 4H-10H. APC coring was terminated due to core disturbance from six successive partial strokes and overpull (Cores 5H-10H). XCB cores 11X to 41X (91.7-377.7 mbsf) were cut with 38.9% recovery. WSTP temperature probes were deployed after Cores 14X, 17X, 22X, and 27X. XCB coring was terminated when core recovery became unsatisfactory (13%) in the last nine cores. The hole was filled with 130 bbl of mud and the bit cleared the rotary table at 2245 on 25 March.

Hole 1007C
The ship was moved 20 m to the southwest, and a RCB BHA similar to that used at the previous sites was run to the seafloor, which was tagged with the bit at 647.3 mbsl. Hole 1007C was initiated at 0340 hr on 25 March and drilled from 0 to 302.0 mbsf. RCB cores were cut from 302.0 to 1235.4 mbsf with 53.9% recovery. Three wiper trips were required to condition the hole while RCB coring. A short trip was made from 504 to 357 mbsf without incident. Below 750 mbsf, overpull and drag increased from 10,000 to 20,000 lb, requiring a second short trip. A third wiper trip was made at 821 mbsf, with 80,000 lb overpull noted at 277 mbsf. The drill string had to be back-reamed out of the hole through numerous tight spots to 78.4 mbsf. No ledges or fill were noted in running pipe back to the bottom of the hole and RCB coring continued from 821.4 to 1235.4 mbsf (total depth). Coring was terminated after penetrating the base of the Neogene. A gel mud sweep was circulated, and a conditioning trip for logs was made to 76.2 mbsf, with overpull at 171, 159, and 102 mbsf (50,000 lb). The bit had to be reamed back to bottom from 1141.0 to 1235.4 mbsf. The bit was released and the pipe was positioned at 111.5 mbsf. A quad-combo log was run to 1158 mbsf. The FMS log had to be canceled because of large hole diameter; therefore, a VSP log was run to 1093 mbsf. The VSP tool could not be pulled back into the bit with 4000 lb overpull. As at Site 1005, the Kinley crimper and cutter were used to cut the logging line, and the VSP tool was retrieved. The ship was under way to Site 1008 at 2018 hr on 3 April.

Sites 1008/1009 Operations Report

166 Table of Contents

Publications Home

ODP Home